Hello, I would like some help with this problem,
Sally,Sarah, and sue worked all summer to save money for school clothes.Sarah saved $50 more than twice as much as Sally,and sue saved $100 more than 1/2 as much as Sally.The total amount saved by the three girls was $
1550. How much did each girl save?
Thank you.

Answer by stanbon(75887)

Sally,Sarah, and sue worked all summer to save money for school clothes.
Sarah(a) saved $50 more than twice as much as Sally(b)
Sue(c) saved $100 more than 1/2 as much as Sally.
The total amount saved by the three girls was $1550.
----------------------------------------------
How much did each girl save?
Equations:
a + b + c = 1550
a = 2b + 50
c = (1/2)b + 100
-----
Substitute for "a" and "c" and solve for "b":
(2b+50) + b + (1/2)b + 100 = 1550
---
(7/2)b + 150 = 1550
(7/2)b = 1400
b = (2/7)*1400
b = $400 (Sally's share)
a = 2b+50 = $850 (Sarah's share)
c = (1/2)b+100 = $300 (Sue's share)
===================
Cheers,
Stan H.
===================